List the domain for the function {(-2,3), (-1,5), (0,7), (1,9)}.
List the domain for the function f(x) = x2 - 4
Evaluate the function g(x) = x2 - 5x + 2 for x = 1.
For the function f(x) = 3x + 2, find f(x+h).

Answer by stanbon(48535)

List the domain for the function {(-2,3), (-1,5), (0,7), (1,9)}.
Domain is the set of "x" values, i.e. -2,-1,0,1
---------------------------
List the domain for the function f(x) = x^2 - 4
"x" can take any value so the domain is All Real Numbers.
----------------------------
Evaluate the function g(x) = x^2 - 5x + 2 for x = 1.
g(1)=1^2-5(1)+2
=1-5+2
=-2
---------------------------

For the function f(x) = 3x + 2, find f(x+h).
f(x+h)= 3(x+h)+2
=3x+3h+2
-----------------
Cheers,
Stan H.